Xavier’s Yellow Zippers Corporation (XYZ Corp) sold 2,200 zippers in 2020 for $6 apiece. Xavier buys the zippers for $2 each and it costs $1 to paint each zipper yellow and add a snazzy logo. Xavier paid $200 a month to rent a sales booth at his local outdoor market, along with a sales commission of $1 per zipper sold (his sister made all the sales on commission). Complete XYZ Corp’s contribution income statement.

b) Xavier is wondering what his break-even point is. Calculate both the break-even revenue and sales volume.

c) Next year, Xavier is planning to sell 3,000 zippers. What will be his margin of safety?

d) Prepare a Cost-Volume-Profit graph for XYZ Corp. Label the axes, lines, loss and profit areas, and break-even point clearly on your graph
Answer #1

a) Contribution Income statement:

Contribution Income Statement Sales Revenue 2200*6 13200 Less: variable costs Cost of zipper 2200*2 -4400 Painting cost 2200*

b) Breakeven point is the point at which there is zero profit and zero loss. It is the point at which contribution = fixed costs. In our example, Contribution per unit = Selling Price - Cost of zipper - cost of painting - sales commission

contribution per unit = 6-2-1-1 = 2

We need total contribution = fixed cost = 2400.

Therefore no. of units to be sold = 2400/2 = 1200 units.

At 1200 units. total contribution = 1200*2 = 2400

Profit/Loss = Contribution - Fixed Cost = 2400-2400 = 0

Break even sales revenue = 1200*6 = 7200.

c) Margin of safety = Actual sales - Breakeven Sales.

At 3000 units, Total actual sales = 3000*6 = 18000. Breakeven sales = 7200.

Margin of safety = 18000-7200 = 10800.

d)

Cost-Volume-Profit Graph of XYZ Corp.

motor Sales 1000 Total Costs PROFIT 2000 80001 70001 Breakeven point 6000 to s 3 5000! LOOD S507 30vot 2000+ Fixed cost TO 00The steps are very simple:

Plot the line of fixed cost horizontal to x-axis becuase it doesnt change with the sales volume.

Next, plot the breakeven point at corresponding breakeven revenue and volume.

Draw the cost line starting from the fixed cost on y axis and passing through breakeven point. Cost is never zero because of fixed cost. That is why cost doesnt begin from zero point.

Draw the sales line starting from zero and passing through breakeven point. When zero units are sold, zero revenue is earned. This is why sales line begins from zero point.

The main point of calculating breakeven point before drawing the graph is to figure out at which point these two lines are going to intersect. That is why both these lines should pass through breakeven point.